If y varies inversely proportionate as x, and y= 3 when x= 4 find x when y=6
Answer by Boreal(15235)

y=k/x
3=k/4
k=12
next part use that k value
6=12/x
6x=12
x=2
does this make sense?
as y doubles, x should be cut in half. It is.